A father is now 6 times as old as his son. In two years, the father will be 5 times as old as his son. Find the present age of the father.
Answer by JulietG(1812)

F = 6S
F+2 = 5(S+2)
Distribute
F+2 = 5S + 10
Subtract 2 from each side
F = 5S +8
Substitute that value of F for the F in the top equation.
F = 6S
(5S+8) = 6S
Subtract 5S from each side
8 = S
If the son is 8, then the father is 48
In two years, the son will be 10 and the father will be 50.